An electric heater supplies 17.2 joules of energy to 27.8-g block of Cu originally at 17.3 °C. Compute the final temperature (in °C).


°C
1
Expert's answer
2016-11-26T13:58:11-0500
The final temperature of the block T = T0 + H/C*m, where T0 - initial temperature of the block (K), H - amount of heat transferred to the block (J), C- specific heat of copper (J/gK), m – mass of the block (g).
The specific heat of copper is 0.385 J/gK (https://en.wikipedia.org/wiki/Heat_capacity).

To convert °C to K we have to add 273.15.
Do the calculation:
T = (17.3+273.15) K + 17.2 J/ (0.385 J/gK * 27.8 g) = 292.06 K or (292.06-273.15) = 18.9 °C.

Answer:
The final temperature of copper block is 18.9 °C.